Find the frequency of AabbCcDdee if parents are AabbCCddEe and AabbccDdee. Options are : a) 0.78%b) 12.5%b) 25%c) 50%Please explain me the solution.

Grade:12th pass

2 Answers

Arun
6 years ago
Dear Amisha
 
Since the inheritance of each gene is independent, So in order to solve this problem, each gene is analysed separately.
 
1. Aa x Aa = 1/4 AA; 1/2 Aa and 1/4 aa. This means 1/2 offsprings have Aa.
2.      bb x bb = all bb = 1.  
3.      CC x cc = all Cc = 1
4. Dd x dd=  1/2​ Dd, 1/2 dd. This means 1/2 offsprings have Dd.
5.      Ee x ee =  1/2​ Ee, 1/2 ee. This means 1/2 offsprings have ee.
Phenotypic proportion that look like AabbCcDdee
 1/2 x 1 x 1 x 1/2 x 1/2 = 1/8.
Hence in% = 100/8 = 12.5%
 
Regards
Arun(askIITians forum expert)
ankit singh
3 years ago
According to the law of independent inheritance, the alleles are inherited independently. Hence, in the given question, each gene is analysed separately.
 
1. Aa x Aa = 1/4 AA, 1/2 Aa and 1/4 aa. Hence, 1/2 or 50% offsprings will have Aa.
2. Similarly, bb x bb = all bb. Hence, all offsprings will have bb.  
3. cc x CC = all Cc. Hence, all ossprings will have Cc. 
4. Dd x dd =  1/2​ Dd and 1/2 dd. Hence, 1/2 or 50% offsprings will have Dd.
5. Ee x ee =  1/2​ Ee and 1/2 ee. Hence, 1/2 or 50% offsprings will have ee.
Hence, frequency of AabbCcDdee will be 1/2 x 1 x 1 x 1/2 x 1/2 = 1/8, which is equal to 12.5%.
Thus, the correct answer is '12.5%.